OpenStudy (anonymous):

How to you factor 8x^3 + 27

OpenStudy (anonymous):

\[8x ^{3}+27\]

OpenStudy (jdoe0001):

\(\bf 8x^3 + 27\qquad notice \begin{cases} 8\to 2\cdot 2\cdot 2\to {\color{brown}{ 2^3}}\\ \bf 27\to 3\to 3\to 3\to {\color{brown}{ 3^3}} \end{cases}\qquad thus \\ \quad \\ 8x^3 + 27\implies 2^3x^3+3^3\implies (2x)^3+3^3 \\ \quad \\ recall\implies {\color{brown}{ a^3+b^3\implies (a+b)(a^2-ab+b^2)}}\)